If the interest rate per year is 16 and the compounding occurs every quarter, then what is the interest rate per compounding period?


1
Expert's answer
2021-01-19T04:31:31-0500

R = Rate of Interest per year (16% = 0.16)

m = times (quarters )

Interest Rate (R)

is the nominal interest rate or "stated rate" in percent. r = R/100

Compounding Periods (m)

is the number of times compounding will occur during a period.

Periodic Interest Rate (P)

This is the rate per compounding period, such as per month when your period is year and compounding is 12 times per year.

Periodic Interest Rate (P) = "\large\frac{Annual Rate (R)}{Compounding (m)}" = "\large\frac{0.16}{4} = 0.04 = 4\%"